twice the smallest of three consecutive odd intergers is seven more
than the largest.
find the integers

Answer by checkley77(12844)

Let x, x+2 & x+4 be the 3 odd imtegers.
2x=(x+4)+7
2x=x+11
2x-1=11
x=11 ans. for the smallest.
11+4=15 ans. for the largest.
Proof:
2*11=15+7
22=22